What is to depreciation life for a project?

Give a clear example of it!
Answer #1

Depreciation is charge of asset over the  useful life of asset or project .There are many methods to calculate depreciation like straight line method,Double Diminishing balance method,Sum of years digit  method ,for tax purpose( MACRS ).

For Example

Purchase cost $ 100000

life 5 year

In straight line method, constant amount is charge over life of asset.

Depreciation = 100000/5= $20000

In Double declining balance method ,Depreciation expense is higher in earlier years .It gradually decrease over time.

Depreciation rate = 1/5*2=20%*2 =40%

Year opening balance depreciation expense closing balance

1 100000 40000 60000

2 60000 24000 36000

3 36000 14400 21600

4 21600 8640 12960

5 12960 5184 7776